OpenStudy (anonymous):

p(t)=32(1-e^-kt)+64 represents the expected score for a student who studies t hours for a test. k=.2 scores must be integers if the student does not study, what score can he expect?

OpenStudy (disqin):

If he doesn't study then t=0. then according to your equation. \[p(t)=32(1-e^{-2\mathrm x 0})+64\] \[\implies p(t)=32(1-0)+64\] \[\implies p(t)=32+64=98\] Wow! Is this question a joke... lol... How one can get that much without studying.

OpenStudy (anonymous):

without studying the highest grade he can get is a 64